# Sustainable Exploitation Equilibria for Dynamic Games

### v2

@@ Abstract
We introduce the Sustainable Exploitation Equilibrium (SEE), a refinement of Markov Perfect Equilibrium (MPE) for dynamic games with an exploiter–exploitee structure. SEE imposes renegotiation-proof exploiter-optimal selection on the set of rationally viable stationary Markov equilibria, where viability follows from sequential rationality when exiting a sustainability set entails catastrophic losses. Unlike MPE, SEE rules out equilibria in which the exploiter optimally drives the state to collapse despite positive continuation payoffs. The exploitee cannot exit, but retains a strategic effort margin affecting dynamics and payoffs. We establish existence under standard conditions, and the refinement is illustrated in a hegemon–client model of foreign politics.
@@ 1 Introduction
Exploitation arises when one agent extracts value from another whose continued functioning is essential for the relationship itself. In such dynamic environments, standard equilibrium concepts such as Markov Perfect Equilibrium (MPE) may admit outcomes in which the exploiter optimally extracts too aggressively, driving the state into collapse despite positive continuation payoffs. These equilibria rely on the exploiter being indifferent to catastrophic breakdown once it occurs. To address this, this paper introduces the Sustainable Exploitation Equilibrium (SEE), a refinement of stationary Markov Perfect Equilibrium that rules out such collapse outcomes by deriving viability from rationality. When exiting a sustainability region entails sufficiently large losses for the exploiter—and when a feasible safe action exists that preserves the relationship—any equilibrium placing positive probability on collapse is sequentially irrational. Under these conditions, restricting attention to viable stationary equilibria is without loss.
SEE then further refines the set of rationally viable stationary MPE by imposing renegotiation-proofness and exploiter-optimal selection. SEE is motivated by settings where one party cannot exit but remains responsive, such as in hegemon–client dynamics, as the exploitee retains a strategic effort margin that shapes both payoffs and dynamics. In this way, SEE formalizes the idea of “sustainable exploitation”: the exploiter extracts as much as possible while ensuring the long-run survival and productivity of the exploitee. This rational-viability foundation distinguishes SEE from equilibrium selection rules that impose sustainability as an external constraint.
@@ 2 Related Literature
SEE builds on the refinement tradition in dynamic games. (MaskinTirole2001a) develop the theory of Markov perfect equilibrium; (Bernheim1987) introduce coalition-proof Nash equilibrium; (Selten1975) trembling-hand perfection; and (FarrellMaskin1989) renegotiation-proofness in repeated games.
@@
The viability constraint draws from (Aubin1991) Viability Theory and constrained Markov decision processes (Altman1999), ensuring state trajectories remain within sustainability sets. This connects to the resource economics literature, e.g. (Dasgupta1979), (Chichilnisky1997), and (Dockner2000), where long-run viability is central to dynamic exploitation of resources.
@@
SEE also relates to political economy models of extraction and state capacity ((AcemogluRobinson2006); (BesleyPersson2011)) and to relational contracting without full commitment ((ThomasWorrall1988); (Levin2003)). In these contexts, weaker parties cannot exit but retain effort margins that shape dynamics. SEE formalizes the logic of “restrained exploitation” within a Markovian framework.
@@ 3 Model and Definition
We interpret transitions outside a sustainability set $V\subset S$ as catastrophic breakdown for the exploiter, in the sense that sufficiently small reductions in extraction can strictly improve continuation payoffs by preserving the relationship. Under this interpretation, equilibrium behavior that places positive probability on exiting $V$ is incompatible with sequential rationality when collapse entails sufficiently large losses.
@@ A Sustainable Exploitation Equilibrium is any stationary MPE $\sigma^{*}\in\mathcal{E}$ satisfying:
    Renegotiation-proofness: At any state $s\in V$, there is no alternative viable profile $\tilde{\sigma}\in\mathcal{E}_{V}$ such that:
@@
    with at least one inequality strict.
@@
    Exploiter-optimality: Among all renegotiation-proof elements of $\mathcal{E}_{V}$, $\sigma^{*}$ maximizes the exploiter’s value $W_{X}(\cdot;s)$ at the state $s$ (with ties broken lexicographically).
@@ 4 Relation to Nash and Markov Perfect Equilibrium
SEE is defined as a viable, renegotiation-proof MPE with exploiter-optimal selection. Formally, a stationary Markov profile $(\sigma_{X},\sigma_{E})$ is an SEE if the aforementioned conditions of Definition 3 are satisfied.
It follows immediately that
@@ 5 Existence under rational viability
We derive viability from rationality by interpreting exits from a sustainability set as catastrophic for the exploiter. We interpret the catastrophe penalty $M$ as a reduced-form representation of franchise loss, regime collapse, or irreversible destruction of the exploitee following breakdown of the relationship.
Assumption (Catastrophic collapse and safe action). Payoffs are bounded: $|u^{X}(s,x,e)|\leq\bar{u}<\infty$. Let $V\subset S$ be a nonempty compact set. For each $s\in V$, there exists a safe action $\underline{x}(s)\in X(s)$ such that, under the exploitee’s best response,
@@
Suppose the state space $S$ and action sets $X(s),E(s)$ are compact, payoffs are bounded and continuous, and the transition kernel $Q$ satisfies the Feller property. Interpret $S\setminus V$ as a collapse region generating a catastrophic loss for the exploiter.
Then there exists $M^{\star}<\infty$ such that, for all catastrophe penalties $M\geq M^{\star}$, every stationary Markov Perfect Equilibrium of the penalized game
@@
Sketch.
Under compactness and continuity, stationary Markov Perfect Equilibria exist. Fix any $s\in V$. If an equilibrium action induces $Q(S\setminus V\mid s,x,e)>0$, the exploiter’s expected one-period payoff is at most $\bar{u}-Mp$ for some $p>0$. Under the safe action $\underline{x}(s)$, the expected penalty is zero and the payoff is at least $\bar{u}$. For $M$ sufficiently large, the safe action strictly dominates any action with positive exit probability. By continuity and compactness, a finite threshold $M^{\star}$ exists. Hence any equilibrium must be viable. Existence of renegotiation-proof and exploiter-optimal selections follows by compactness and continuity. ∎
@@ 6.4 Follower’s best response
An interior first-order condition is
$\phi^{\prime}(e^{*})\;+\;\delta\,W^{C}_{s}\!\big(s^{\prime}\big)\,f_{e}(s,e^{*})\;=\;0,\qquad s^{\prime}\equiv f(s,e^{*})-h(x).$
Under standard concavity/regularity, the implicit function theorem yields $\partial e^{*}/\partial x$.
@@ 6.5 Leader’s SEE problem
subject to viability $x\leq f(s,e^{*}(s,x))-s_{\min}$. Let $\mu\geq 0$ be the multiplier on viability. The Lagrangian is
$\mathcal{L}(x;s,\mu)\;=\;\pi(x)-k(s,e^{*})+\delta V^{H}(s^{\prime})\;+\;\mu\big(f(s,e^{*})-h(x)-s_{\min}\big),$
where $e^{*}\equiv e^{*}(s,x)$ and $s^{\prime}\equiv f(s,e^{*})-h(x)$.
The (interior) first-order condition with envelope terms from $e^{*}(s,x)$ is
$\pi^{\prime}(x)\;-\;\delta V^{H}_{s}(s^{\prime})\,\big(h^{\prime}(x)-f_{e}(s,e^{*})\,\partial e^{*}/\partial x\big)\;-\;\mu\,\big(h^{\prime}(x)-f_{e}(s,e^{*})\,\partial e^{*}/\partial x\big)\;=\;0,$
with complementary slackness $\mu\big(f(s,e^{*})-h(x)-s_{\min}\big)=0$.
Interpretation.
The hegemon equates marginal extraction benefits $\pi^{\prime}(x)$ to the discounted shadow cost of lost future capacity $h^{\prime}(x)$ net of the induced effort response $f_{e}\partial e^{*}/\partial x$. When viability binds, the shadow price $\mu>0$ further restrains extraction.
6.6 Stationary SEE and regimes
A stationary SEE consists of $(x^{*}(s),e^{*}(s,x))$ and an invariant distribution supported on $V=\{s\geq s_{\min}\}$. Under rational viability, two canonical regimes arise:
  • Interior (viability slack, $\mu=0$). If the unconstrained fixed point $s^{*}$ of $s^{\prime}=f(s,e^{*}(s,x^{*}))-h(x^{*})$ satisfies $s^{*}>s_{\min}$, the FOCs imply a generalized “golden rule” where effective regeneration $f_{s}+f_{e}\,\partial e^{*}/\partial s$ balances discounting (details depend on primitives).
  • Boundary (viability binds, $\mu>0$). If the unconstrained $s^{*}$ would fall below $s_{\min}$, SEE selects $s^{*}=s_{\min}$ and $x^{*}=f(s_{\min},e^{*}(s_{\min},x^{*}))-s_{\min}$, with $\mu$ set to satisfy the FOC.
Additional Remark on Outside Option
One might alternatively assume that the exploitee $E$ has an outside option $U^{out}_{E}(s)$, in which case viability must be supplemented by an individual rationality (IR) constraint: $W^{E}(s)\geq U^{out}_{E}(s)$ for all $s\in V$. This yields an even tighter refinement, as it restricts feasible equilibria to those in which the exploitee weakly prefers participation over exit.
@@ 7 Conclusion
With a non-exiting but strategically responsive exploitee, SEE refines Markov Perfect Equilibrium by deriving viability from rationality and imposing renegotiation-proof exploiter-optimal selection. Unlike MPE, SEE eliminates equilibria in which the exploiter optimally drives the state to collapse despite positive continuation payoffs. In equilibrium, SEE yields either an interior “golden-rule”–type extraction path or, when unconstrained extraction would cause collapse, a boundary outcome in which the state is pinned at the viability threshold. The framework applies broadly to settings in which one party cannot exit but retains an effort margin—such as colonial relationships, labor arrangements, and locked-in suppliers—providing a compact refinement for analyzing persistent exploitative interactions in economics and political economy.

### v3

@@ Abstract
We study dynamic relationships in which one party extracts current surplus in ways that degrade the future state, while the counterparty cannot exit but adjusts effort in response. Standard stationary Markov equilibria may sustain collapse paths in which short-run extraction dominates strictly positive continuation gains. We introduce the Sustainable Exploitation Equilibrium (SEE), a refinement for dynamic games with irreversible failure modeled as an absorbing boundary that eliminates continuation value. When a survival-preserving action exists and failure destroys future surplus, equilibria assigning positive probability to collapse are sequentially irrational. Equilibrium analysis can therefore be restricted, without loss, to continuation-preserving stationary Markov equilibria. Within this restricted domain, viable renegotiation-proofness becomes structural: because failure truncates future surplus, any jointly improving survival-preserving deviation is credible prior to collapse. SEE selects the viable, renegotiation-stable equilibrium that maximizes the exploiter’s value. Existence is established under standard conditions, and the refinement is illustrated in a hegemon–client setting.
@@ 1 Introduction
Exploitation arises when one party extracts value from another whose continued functioning is necessary for the relationship itself. In many economic and political environments, aggressive extraction can erode the very state that sustains the interaction—whether by depleting productive capacity, destabilizing institutions, or undermining cooperation. A central strategic tension therefore arises: the actor with extraction power benefits from pushing the system harder today, but doing so risks destroying the future surplus the relationship generates.
Standard equilibrium concepts such as Markov Perfect Equilibrium (MPE) do not fully resolve this tension. Because equilibria are evaluated over all admissible state transitions, they may sustain collapse paths in which short-run extraction dominates strictly positive continuation values. This paper introduces the Sustainable Exploitation Equilibrium (SEE), a refinement of stationary MPE for environments with irreversible failure.
Why the refinement is necessary under irreversible failure.
In dynamic environments without absorbing collapse, continuation payoffs are defined on a common domain across strategies. Irreversible failure changes this structure: once the state exits the survival region, the continuation problem collapses. Standard equilibrium notions evaluate strategies without restricting the admissible continuation domain, thereby permitting profiles that rely on the destruction of future surplus along the equilibrium path. SEE makes this restriction explicit by requiring that equilibrium paths preserve the continuation domain before selection among them. The refinement therefore reflects a consistency requirement imposed by irreversible failure rather than an additional constraint.
@@ 2 Related Literature
SEE builds on the refinement tradition in dynamic games. (Maskin and Tirole, 2001) develop the theory of Markov perfect equilibrium; (Bernheim et al., 1987) introduce coalition-proof Nash equilibrium; (Selten, 1975) trembling-hand perfection; and (Farrell and Maskin, 1989) renegotiation-proofness in repeated games.
@@
The viability constraint draws from (Aubin, 1991) Viability Theory and constrained Markov decision processes (Altman, 1999), ensuring state trajectories remain within sustainability sets. This connects to the resource economics literature, e.g. (Dasgupta, 1979), (Chichilnisky, 1997), and (Dockner et al., 2000), where long-run viability is central to dynamic exploitation of resources.
@@
SEE also relates to political economy models of extraction and state capacity ((Acemoglu and Robinson, 2006); (Besley and Persson, 2011)) and to relational contracting without full commitment ((Thomas and Worrall, 1988); (Levin, 2003)). In these contexts, weaker parties cannot exit but retain effort margins that shape dynamics. SEE formalizes the logic of restrained exploitation within a Markovian framework.
Comparison.
In contrast to the above, this paper studies environments in which exit from a viability region constitutes irreversible failure that eliminates continuation value. Standard refinements select among equilibria defined on a fixed continuation domain. SEE instead restricts the admissible domain by requiring equilibrium paths to preserve the survival region before equilibrium selection. Viability and renegotiation-proofness therefore follow from sequential rationality under irreversible failure.
@@ 3 Model and Definition
We interpret transitions outside a sustainability set $V\subset S$ as irreversible failure: absorption into a collapse region that eliminates continuation value. Under this interpretation, equilibrium behavior that places positive probability on exiting $V$ is incompatible with sequential rationality when failure entails sufficiently large losses and a survival-preserving action exists.
@@ A Sustainable Exploitation Equilibrium is any stationary MPE $\sigma^{*}\in\mathcal{E}$ satisfying:
    Renegotiation-proofness: At any state $s\in V$, there is no alternative viable profile $\tilde{\sigma}\in\mathcal{E}_{V}$ such that
@@
    with at least one inequality strict. Renegotiation is modeled in reduced form: because irreversible failure eliminates continuation value, any jointly improving survival-preserving deviation is credible prior to collapse. Renegotiation-proofness therefore acts as a stability requirement on continuation paths within the survival domain.
@@
    Exploiter-optimality: Among all renegotiation-proof elements of $\mathcal{E}_{V}$, $\sigma^{*}$ maximizes the exploiter’s value $W_{X}(\cdot;s)$ at state $s$ (ties broken lexicographically).
@@
Multiplicity and uniqueness.
The viability restriction removes collapse-inducing equilibria, while renegotiation-proofness eliminates survival-preserving equilibria that leave joint surplus unexploited. SEE therefore reduces the equilibrium correspondence relative to MPE and may yield uniqueness, although multiplicity can persist when several survival-preserving equilibria generate identical joint values.
@@ 4 Relation to Nash and Markov Perfect Equilibrium
SEE is defined as a viable, renegotiation-proof MPE with exploiter-optimal selection. It follows immediately that
@@
SEE is strictly stronger than renegotiation-proof MPE.
Renegotiation-proofness selects among equilibria within a given continuation domain but does not restrict that domain itself. With an absorbing collapse state, equilibria may drive the system to failure when short-run gains exceed continuation values and yet remain renegotiation-proof because collapse eliminates future surplus. SEE excludes such outcomes by restricting attention to survival-preserving paths, thereby refining both equilibrium selection and the admissible continuation domain. Similar observations apply to coalition-proof refinements, which eliminate profitable coalitional deviations but likewise evaluate equilibria on a fixed continuation domain.
@@ 5 Existence under rational viability
We derive viability from rationality by interpreting exits from a sustainability set as irreversible failure.
Economic interpretation of irreversible failure.
Irreversible failure is modeled as an absorbing boundary that eliminates continuation value. Economically this corresponds to events such as sovereign default, political collapse, financial liquidation, or irreversible depletion of productive capacity. When the discounted continuation loss exceeds the maximal short-run gain from over-extraction, survival-preserving behavior follows from sequential rationality.
Assumption (Irreversible failure and safe action). Payoffs are bounded: $|u^{X}(s,x,e)|\leq\bar{u}<\infty$. Let $V\subset S$ be a nonempty compact set. For each $s\in V$, there exists a safe action $\underline{x}(s)\in X(s)$ such that, under the exploitee’s best response,
@@
Assumption (Uniform exit risk for non-survival-preserving actions). For each $s\in V$, let
$\mathcal{X}_{s}^{\,N}\;=\;\{x\in X(s):Q(V\mid s,x,\sigma_{E}(s,x))<1\}$
denote the set of actions that are not survival-preserving (given the exploitee’s best response). Assume that the minimal probability of exiting $V$ over all such actions is bounded away from zero:
$p^{\star}\;=\;\inf_{s\in V}\inf_{x\in\mathcal{X}_{s}^{\,N}}\big(1-Q(V\mid s,x,\sigma_{E}(s,x))\big)\;>\;0.$
@@
Suppose the state space $S$ and action sets $X(s),E(s)$ are compact, payoffs are bounded and continuous, and the transition kernel $Q$ satisfies the Feller property. Interpret $S\setminus V$ as an absorbing failure region that eliminates continuation value.
Then there exists $M^{\star}<\infty$ such that, for all failure penalties $M\geq M^{\star}$, every stationary Markov Perfect Equilibrium of the penalized game
@@
Proof.
Under compactness and continuity, stationary Markov Perfect Equilibria exist. Fix any $s\in V$. If an equilibrium action induces $Q(S\setminus V\mid s,x,e)>0$, the exploiter’s expected one-period payoff is at most $\bar{u}-Mp$ for some $p^{*}>0$. Under the safe action $\underline{x}(s)$, the expected penalty is zero and the payoff is at least $\bar{u}$. For $M$ sufficiently large, the safe action strictly dominates any action with positive exit probability. By continuity and compactness, a finite threshold $M^{\star}$ exists. Hence any equilibrium must be viable. Existence of renegotiation-proof and exploiter-optimal selections follows by compactness and continuity. ∎
Finite collapse threshold.
Because payoffs are bounded and safe actions exist at every state in the survival domain, there exists a finite failure threshold $M^{*}$ such that any action inducing a positive probability of exit is strictly dominated once $M\geq M^{*}$. The threshold depends on (i) the maximal one-period gain from over-extraction, (ii) the minimal probability of exit under non-safe actions, and (iii) the discount factor. In particular, higher exit risk under over-extraction lowers the required threshold, while greater patience reduces the attractiveness of failure paths.
@@ 6.4 Follower’s best response
Remark. If $W_{C}$ is differentiable, the interior best response satisfies
$\phi^{\prime}(e^{*})+\delta W_{C}^{\prime}(s^{\prime})f_{e}(s,e^{*})=0.$
@@ 6.5 Leader’s SEE problem
subject to viability $x\leq f(s,e^{*}(s,x))-s_{\min}$. Let $\mu\geq 0$ be the multiplier on viability.
Additional Implementative Remarks
Robustness to small outside options.
The no-exit assumption sharpens the exploitation structure but is not essential. Suppose the exploitee has an outside option yielding value at most $\varepsilon>0$ relative to continuation within the survival domain. For sufficiently small $\varepsilon$, survival-preserving stationary Markov equilibria and the renegotiation-stable selection vary continuously. When collapse losses for the exploiter exceed the maximal short-run gain from over-extraction, the viability restriction continues to bind and SEE converges to the baseline no-exit refinement as $\varepsilon\to 0$.
Timing and simultaneity.
The baseline adopts sequential timing in which extraction precedes effort adjustment, but the refinement does not depend on this structure. Under simultaneous moves, stationary Markov equilibria remain well-defined, and the survival-domain restriction continues to exclude collapse-inducing profiles. The mechanism therefore reflects the truncation of continuation value at failure rather than the timing protocol.
Empirical operationalization.
The state $s_{t}$ can be proxied by observable fiscal or administrative capacity indices (e.g. World Bank Governance Indicators), extraction $x_{t}$ by documented concession or compliance flows, and $s_{\min}$ by sovereign default or regime-transition events observable in panel data. Structural parameters governing persistence and responsiveness can be calibrated using standard approaches in state-capacity models (e.g. Besley and Persson (2011)). The viability threshold is disciplined by the present value of foregone extraction rents following collapse.
@@ 7 Conclusion
Dynamic relationships often involve asymmetric control over current surplus and future viability. When extraction degrades the state sustaining the interaction, standard equilibrium concepts may permit collapse paths even when continuation values remain positive.
We introduce the Sustainable Exploitation Equilibrium (SEE), a refinement of stationary Markov Perfect Equilibrium for environments with irreversible failure. Because failure eliminates continuation value, survival-preserving deviations are credible prior to collapse, making renegotiation-proofness endogenous and restricting equilibrium analysis to viable stationary equilibria. If unconstrained extraction preserves viability, SEE yields an interior outcome; if it would induce failure, equilibrium instead converges to the viability boundary—the maximal level of exploitation consistent with survival.
SEE thus provides a compact framework for analyzing persistent exploitative relationships while clarifying how irreversible failure disciplines incentives without eliminating pressures toward maximal sustainable exploitation.